Given the equations below, calculate the equilibrium constant for the reaction:

2 A(g) + B(g) ⇌ A2B(g)3 A(g) + 3 C(g) ⇌ 3 AC(g)        K = 8.02

B(g) + 2 AC(g) ⇌ 2 AB + 2 C(g)                                    K = 25

A2B(g) ⇌A(g) + AB(g)                                                   K = 10
Answer #1

3 A(g) + 3 C(g) ⇌ 3 AC(g)        K = 8.02

B(g) + 2 AC(g) ⇌ 2 AB + 2 C(g)                                    K = 25

A2B(g) ⇌A(g) + AB(g)      K =10

The required reaction is:

2 A(g) + B(g) ⇌ A2B(g)

This can be written in terms of given 3 reactions

This reaction = (1/3)*reaction1 + (1/2)*reaction 2 -1*reaction 3

So,

K = K1^(1/3)*k2^1/2 / k3

= 8.02^1/3 * 25^1/2 / 10

= 2.00 * 5 / 10

= 1

Answer: 1